# Generative thermodynamic computing

### v1

@@ Abstract
In more detail, we consider a model thermodynamic computer, a set of fluctuating nonlinear degrees of freedom coupled by bilinear interactions. The computer’s degrees of freedom evolve according to overdamped Langevin dynamics. This design is inspired by existing hardware that can perform linear algebra Aifer et al. (2024), and by our recent work showing that a nonlinear version of such hardware can function as the thermodynamic version of a neural network Whitelam and Casert (2024). We provide input to the computer to make it display images of digits from the MNIST data set LeCun et al. (1998), and allow these images to degrade by running the dynamics of the computer with its inter-unit couplings set to zero. Such degradation is called noising in the diffusion model literature. As we do so, we compute from the Langevin equation the probability that a computer with hypothetical nonzero couplings would have generated the reverse of this noising trajectory, and we adjust the values of these hypothetical couplings by gradient descent in order to maximize that probability. After running several such noising trajectories, we construct a denoising computer using the trained couplings, and verify that its natural dynamics, starting from noisy initial conditions, leads to the generation of structured MNIST-like digits. Independent dynamical trajectories of the same computer produce a variety of outcomes, some of which were not contained in the training set.
@@
Imagine that we observe a dynamical trajectory of the computer at a series of discrete times, $\omega=\{{\bm{x}}(t_{k})\}_{k=0}^{K}$, where $t_{k}=k\Delta t$. The probability that any step of this trajectory was generated by a thermodynamic computer with parameters ${\bm{\theta}}$ can be calculated from Eq. (1). One way to do so is to consider the dynamical action associated with the Langevin equation, but it is simpler to consider a typical integration scheme, such as the Euler scheme, that is used on a classical digital computer to integrate the equation:
@@
Numerical illustration of generative thermodynamic computing — To illustrate this result we carried out a digital simulation of a thermodynamic computer. We set $J_{2}=J_{4}=10k_{\rm B}T$, and consider a computer with $N_{\rm v}=28^{2}$ visible units and $N_{\rm h}=128$ hidden units. The visible units will be used as a display, and the hidden units used to do computation. The computer has $N_{\rm v}N_{\rm h}$ trainable couplings $J_{ij}$ between visible and hidden units, and $N_{\rm h}(N_{\rm h}-1)/2$ trainable couplings between hidden units. Hidden units have trainable biases $b_{i}$. Visible units have no trainable biases, but during training we impose visible-unit biases $b_{i}\propto P_{i}$, where $P_{i}$ denotes the $i^{\rm th}$ pixel of an MNIST digit (each digit’s pixels were adjusted to have zero mean and unit variance). We display the visible units in a $28\times 28$ grid, matching the presentation of an MNIST digit.
To construct noising trajectories we project an MNIST image onto the visible units, via their non-trainable biases. We set all other biases and couplings to zero, and let the computer come to equilibrium, by simulating Eq. (1) for a sufficiently long time. We then run a dynamical trajectory of time $t_{\rm f}=2.5$, slowly diminishing the intensity of the imposed digit. The result is an image that becomes increasingly noisy, as shown in Fig. 1(a). Positive values of the unit activations $x_{i}$ are shown blue, while negative values are shown white.
@@
In Fig. 2(a) we show three independent trajectories of the thermodynamic computer trained in this way. Trajectories begin from a noisy initial state prepared by bringing the coupling-free thermodynamic computer to equilibrium. In each case, the trained thermodynamic computer gradually transforms noise into structure, without external control, demonstrating its generative capabilities. These results indicate that the computer has internalized representations of the digits and can reproduce them via physical evolution. Independent trajectories yield distinct samples, indicating stochastic generation rather than memory retrieval.
Fig. 2(b) we show the output at time $t=t_{\rm f}$ of 25 independent trajectories generated by the trained thermodynamic computer. The outputs exhibit diversity in style and form, indicating that the system is not simply memorizing inputs but generating configurations that interpolate within the learned distribution. This generative behavior arises purely from the system’s autonomous physical evolution, without external control. Not all outputs are necessarily desirable or clearly interpretable – some display “mode mixing”, in the language of diffusion models – but this simple example is intended as a proof of principle, a demonstration that structure can be generated from noise by physical dynamics alone.
Some insight into the behavior of the computer is provided by Fig. 2(c), which shows 16 of the computer’s hidden units’ learned couplings to the visible layer. Each panel represents the pattern of couplings between one hidden unit and all visible units: blue indicates positive coupling, red negative (either can be inhibitory, depending on the values of the unit activations $x_{i}$). These patterns act as receptive fields that determine how each hidden unit responds to features in the visible layer. Several of the units have learned localized, digit-like structures, suggesting that the computer decomposes input patterns into interpretable visual components. Some hidden units exhibit complementary structure – activating for certain strokes while suppressing others – enabling the system to encode multiple features. Together, these receptive fields determine the energy landscape that guides the computer’s dynamics, allowing the system to transform noise into structured outputs.
@@
Acknowledgments— I thank Isaac Tamblyn for discussions. This work was done at the Molecular Foundry, supported by the Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231, and partly supported by US DOE Office of Science Scientific User Facilities AI/ML project “A digital twin for spatiotemporally resolved experiments”.

### v2

@@ Abstract
In more detail, we consider a model thermodynamic computer, a set of fluctuating nonlinear degrees of freedom coupled by bilinear interactions. The computer’s degrees of freedom evolve according to overdamped Langevin dynamics. This design is inspired by existing hardware that can perform linear algebra Aifer et al. (2024), and by our recent work showing that a nonlinear version of such hardware can function as the thermodynamic version of a neural network Whitelam and Casert (2024). We provide input to the computer to make it display images of digits from the MNIST data set LeCun et al. (1998), and allow these images to degrade by running the dynamics of the computer with its inter-unit couplings set to zero. Such degradation is called noising in the diffusion model literature. As we do so, we compute from the Langevin equation the probability that a computer with hypothetical nonzero couplings would have generated the reverse of this noising trajectory, and we adjust the values of these hypothetical couplings by gradient descent in order to maximize that probability. After running several such noising trajectories, we construct a denoising computer using the trained couplings, and verify that its natural dynamics, starting from noisy initial conditions, leads to the generation of structured MNIST-like digits. Independent dynamical trajectories of the same computer produce a variety of outcomes, some of which are not contained in the training set.
@@
Imagine that we observe a dynamical trajectory of the computer at a series of discrete times, $\omega=\{{\bm{x}}(t_{k})\}_{k=0}^{K}$, where $t_{k}=k\Delta t$. The probability that any step of this trajectory was generated by a thermodynamic computer with parameters ${\bm{\theta}}$ can be calculated from the Onsager-Machlup action associated with the Langevin equation Risken and Risken (1996); Cugliandolo and Lecomte (2017). A time-discretized version of this action can be derived by first considering a standard Euler integration scheme for Eq. (1),
@@
Numerical illustration of generative thermodynamic computing — To illustrate this result we carried out a digital simulation of a thermodynamic computer. We set $J_{2}=J_{4}=10k_{\rm B}T$, and consider a computer with $N_{\rm v}=28^{2}$ visible units and $N_{\rm h}=512$ hidden units. The visible units will be used as a display, and the hidden units used to do computation. The computer has $N_{\rm v}N_{\rm h}$ trainable couplings $J_{ij}$ between visible and hidden units, and $N_{\rm h}(N_{\rm h}-1)/2$ trainable couplings between hidden units. Hidden units have trainable biases $b_{i}$. Visible units have no trainable biases, but during training we impose visible-unit biases $b_{i}\propto P_{i}$, where $P_{i}$ denotes the $i^{\rm th}$ pixel of an MNIST digit (each digit’s pixels were adjusted to have zero mean and unit variance). We display the visible units in a $28\times 28$ grid, matching the presentation of an MNIST digit.
To construct noising trajectories we project an MNIST image onto the visible units, via their non-trainable biases. We project part of the same image onto the hidden-unit biases, in order to provide them with some signal, and set all couplings $J_{ij}$ to zero. We then let the computer come to equilibrium, by simulating Eq. (1) for a sufficiently long time. We then run a dynamical trajectory of time $t_{\rm f}=2.5$, slowly diminishing the intensity of the imposed digit. The result is an image that becomes increasingly noisy, as shown in Fig. 1(a). Positive values of the unit activations $x_{i}$ are shown blue, while negative values are shown white.
@@
In Fig. 2(a) we show three independent trajectories of the thermodynamic computer trained in this way. Trajectories begin from a noisy initial state prepared by bringing the coupling-free thermodynamic computer to equilibrium. In each case, the trained thermodynamic computer gradually transforms noise into structure, illustrating its ability to perform noise-to-structure generation. These results indicate that the computer has internalized representations of the digits and can reproduce them via physical evolution. In the case of the digit “1”, the generated image appears inverted. This behavior reflects a breaking of the visible layer’s approximate symmetry under the transformation $x_{i}\to-x_{i}$ (the digits presented to the computer have pixel values that are roughly symmetric about zero).
Fig. 2(b) we show the output at time $t=t_{\rm f}$ of 25 independent trajectories generated by the trained thermodynamic computer. The outputs exhibit diversity in style and form, behavior that is typical of diffusion models trained to approximate, rather than exactly replicate, the data distribution. Not all outputs are necessarily desirable or clearly interpretable – some display “mode mixing”, in the language of diffusion models – but this simple example is intended as a proof of principle, a demonstration that structure can be generated from noise by physical dynamics alone.
@@
Some insight into the behavior of the computer is provided by Fig. 2(c), which shows 16 of the computer’s hidden units’ learned couplings to the visible layer. Each panel represents the pattern of couplings $J_{ij}$ between one hidden unit and all visible units: blue indicates positive coupling, red negative (either can be excitatory or inhibitory, via the interaction $J_{ij}x_{i}x_{j}$, depending on the values of the unit activations $x_{i}$). These patterns act as receptive fields that determine how each hidden unit responds to features in the visible layer. Several of the units have learned localized, digit-like structures, suggesting that the computer decomposes input patterns into interpretable visual components. Some hidden units exhibit complementary structure – activating for certain strokes while suppressing others – enabling the system to encode multiple features. Together, these receptive fields determine the energy landscape that guides the computer’s dynamics, allowing the system to transform noise into structured outputs.
@@
Acknowledgments— I thank Isaac Tamblyn and Adrianne Zhong for discussions. This work was done at the Molecular Foundry, supported by the Office of Science, Office of Basic Energy Sciences, of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231, and partly supported by US DOE Office of Science Scientific User Facilities AI/ML project “A digital twin for spatiotemporally resolved experiments”.